# Training Plug-and-Play Knowledge Modules with Deep Context Distillation

### v3

@@ 2.4 Task Adaptation with Knowledge Extractors
where [.][.][ . ] denotes our combination function. To combine $\theta_{\texttt{KE}}$ and $\theta_{\texttt{KM}}^{D_{i}}$, we use a learnable weighted combination of the two LoRAs applied after the outer product: $[\theta_{\texttt{KM}}^{D_{i}},\theta_{\texttt{KE}}]_{w}=w_{M}A_{D_{i}}B_{D_{i}}^{T}+w_{E}A_{E}B_{E}^{T}$, where $w_{M},w_{E}$ are the scalar weights given to the KM and KE parameters respectively. We learn different combination parameters for every layer where LoRA is applied. We will show in the experiments that, although the KE is trained in tandem with a set of documents during training, it can generalize well to extracting knowledge from an unseen set of documents and corresponding KMs in the held-out set.
@@ Methods
[…]
@@ Results
We report the results for the two base models and the two datasets in Table 1. Results are consistent and show that SDCD performs best across the board in the closed book setting, outperforming both LM, PIT and DDCD. The result hold if we train on Summary, QA, or both, the latter performing the best. In the open book setting, we see that KMs struggle at zero-shot combination with RAG (RAG + KM lines). For Llama-3.1, there are slight signs of forgetting (comparing RAG vs. RAG + KM), which are due to the fact that KMs are never trained with long contexts. Indeed, we see that training with longer concatenated samples addresses this gap, showing consistent gains over RAG. Moreover, we see a similar performance boost when training a specialized KE, which highlights strongly synergistic behavior of RAG and KMs. Comparing RAG + KE and RAG + KM + KE, we see 4.2 & 4.1 Rouge-L gains on NQA and 2.4 & 4.1% Accuracy improvements on QuALITY, for Phi-3 & Llama 3.1 respectively. Lastly, we note that in settings where no synthetically generated data is available, our distillation objective (DDCD) proves to be a better learning signal than next-work prediction (LM).
@@
Finally, we explore if the different components of our DCD loss are useful. In Tab. 4, we show that additionally matching the hidden states does provide benefits in settings where there are no KEs. Indeed, combining both KL and $L_{1}$ losses gives the best performance. When a task-specific KE is available, we see that the KE can mitigate this difference in performance gap. We also experiment with varying the number of synthetic datapoints available per document chunk. In Tab. 5 we show that as we increase the number of synthetically generated data for each chunk, the final performance is increased. Overall, increasing the amount of offline computation in the synthetic generation step, either by generating more data, or from a stronger model, leads to higher performance.
@@ 4 Related Work
Knowledge Injection with Modules Several works have similar goals of KMs. Xiao et al. (2023) introduce “plug-and-play” document modules where an MLP transforms encoded document tokens into soft prompts used across attention layers. Unlike the work in Xiao et al. (2023), we require that the inference compute is fixed and does not scale with the length of the document $D$Zhang et al. (2023a) similarly train knowledge modules in a task-agnostic or task-specific manner, caching the processed document representation for efficient reuse. Amortization-based approaches train a “hypernetwork” to generate lightweight adaptations (e.g., low-rank shifts or prefix-tuning parameters) from a given context. Chen et al. (2024) learn to project context tokens into low-rank parameter shifts in a self-supervised manner, and Tack et al. (2024) encode documents into prefix-tuning weights using a T5-based hypernetwork. These methods train the hypernet on multi-task data, so at inference time it can produce task-specific or document-specific modules in a single forward pass. KMs as described in this paper are trained with gradient descent on single documents independent of any multi-task training. This increases per-document costs but reduces the risk of domain shift. Future work might be devoted to efficient learning of KMs.
@@ Appendix B Ablations
The following ablations are done on a 140 document subset of NQA (100 train, 20 dev, 20 test).

### v4

@@ 2.4 Task Adaptation with Knowledge Extractors
where $[.]$ denotes our combination function. To combine $\theta_{\texttt{KE}}$ and $\theta_{\texttt{KM}}^{D_{i}}$, we use a learnable weighted combination of the two LoRAs applied after the outer product: $\theta_{\texttt{KM}}^{D_{i}},\theta_{\texttt{KE}}]_{w}=w_{M}A_{D_{i}}B_{D_{i}}^{T}+w_{E}A_{E}B_{E}^{T}$, where $w_{M},w_{E}$ are the scalar weights given to the KM and KE parameters respectively. We learn different combination parameters for every layer where LoRA is applied. We will show in the experiments that, although the KE is trained in tandem with a set of documents during training, it can generalize well to extracting knowledge from an unseen set of documents and corresponding KMs in the held-out set.
@@ Methods
In the open book setting, we use RAG and ICL as baselines; for RAG, we split each document into passages of 256 tokens and use SFR-embedding-2 (Meng et al., 2024) to embed passages and questions to perform retrieval of the top-5 relevant passages as measured by the cosine similarity with each question. Similarly to KM, we assume we know the document the questions relate to and therefore we only retrieve passages from that document (we don’t search over the set of all documents in the dataset). We report results of zero-shot RAG (just denoted as RAG) and a fine-tuned version of RAG (RAG + KE), where a KE module (just a LoRA adapter) is fine-tuned on the task with the RAG context. For all methods, KEs are always trained solely on the training documents, and never on the dev/test documents. We experiment with combinations of RAG and KMs, namely RAG + KM (zero-shot), and RAG + KE + KM (KE trained to combine RAG and KM information) to analyze the synergy between RAG and KMs. Technically, ICL for decoder-only models can be considered as a closed book approach if the KV document cache is stored in memory. However, this comes at an excessive storage cost (for Llama3.1 8B, for a document of 60k tokens, it’s $\sim$30Gb). We experiment with KV compression methods such as the recently proposed L2 compress (Devoto et al., 2024b) to analyze tradeoffs between performance and storage cost.
@@ Results
We report the results for the two base models and the two datasets in Table 1. Results are consistent and show that SDCD performs best across the board in the closed book setting, outperforming both LM, PIT and DDCD. The result hold if we train on Summary, QA, or both, the latter performing the best. In the open book setting, we see that KMs struggle at zero-shot combination with RAG (RAG + KM lines). For Llama-3.1, there are slight signs of forgetting (comparing RAG vs. RAG + KM), which are due to the fact that KMs are never trained with long contexts. Indeed, we see that training with longer concatenated samples addresses this gap, showing consistent gains over RAG. Moreover, we see a similar performance boost when training a specialized KE, which highlights strongly synergistic behavior of RAG and KMs. Comparing RAG + KE and RAG + KM + KE, we see 4.2 & 4.1 Rouge-L gains on NQA and 2.4 & 4.1% Accuracy improvements on QuALITY, for Phi-3 & Llama 3.1 respectively. Lastly, we note that in settings where no synthetically generated data is available, our distillation objective (DDCD) proves to be a better learning signal than next-work prediction (LM). A more thorough analysis of this observation is discussed in B.2.
@@
Next, we explore if the different components of our DCD loss are useful. In Tab. 4, we show that additionally matching the hidden states does provide benefits in settings where there are no KEs. Indeed, combining both KL and $L_{1}$ losses gives the best performance. When a task-specific KE is available, we see that the KE can mitigate this difference in performance gap. We also experiment with varying the number of synthetic datapoints available per document chunk. In Tab. 5 we show that as we increase the number of synthetically generated data for each chunk, the final performance is increased. Overall, increasing the amount of offline computation in the synthetic generation step, either by generating more data, or from a stronger model, leads to higher performance.
Lastly, we investigate whether applying specialized KMs leads to forgetting of general knowledge. We evaluate Llama 3.1 (8B) augmented with a random Knowledge Module on the multiple-choice MMLU dataset (Hendrycks et al., 2021), used to measure the model’s knowledge on a wide variety of topics. We find that performance (accuracy) is only negligibly affected, dropping 1% from 63.9% to 62.9%. More details can be found in B.3.
@@ 4 Related Work
Knowledge Injection with Modules Several works have similar goals of KMs. Xiao et al. (2023) introduce “plug-and-play” document modules where an MLP transforms encoded document tokens into soft prompts used across attention layers. Unlike the work in Xiao et al. (2023), we require that the inference compute is fixed and does not scale with the length of the document DDZhang et al. (2023a) similarly train knowledge modules in a task-agnostic or task-specific manner, caching the processed document representation for efficient reuse. Amortization-based approaches train a “hypernetwork” to generate lightweight adaptations (e.g., low-rank shifts or prefix-tuning parameters) from a given context. Chen et al. (2024) learn to project context tokens into low-rank parameter shifts in a self-supervised manner, and Tack et al. (2024) encode documents into prefix-tuning weights using a T5-based hypernetwork. These methods train the hypernet on multi-task data, so at inference time it can produce task-specific or document-specific modules in a single forward pass. KMs as described in this paper are trained with gradient descent on single documents independent of any multi-task training. This increases per-document costs but reduces the risk of domain shift. Future work might be devoted to efficient learning of KMs.
@@ Appendix B Ablations
B.1 Impact of different losses
The following ablation is done on a 140 document subset of NQA (100 train, 20 dev, 20 test).
@@ B.1 Impact of different losses
B.2 Impact of Document DCD over standard next-token prediction
In this section we investigate why Document DCD outperforms the standard language modeling objective. We hypothetise that this is due to the richer signal that the distillation objective provides. Indeed, both the hidden state matching loss, as well as the KL objective, yield more information to the learner than the standard next-token prediction objective. This result aligns with the general finding that distillation (teacher model matching) is more data-efficient than NTP (data matching) in the literature (He et al., 2022; Minixhofer et al., 2025).
To prove this, we run the following experiment: We progressively strip components of DDCD, making it closer to the LM baseline; we remove the l1 loss on the hidden states, and progressively lower the temperature when computing the teacher probabilities $p_{teacher}=\text{softmax}(\frac{logits}{\tau})$, until The target distribution becomes one-hot, effectively mimicking the properties of next-token-prediction loss.
Model Rouge-L
DDCD 21.2
DDCD without hidden state loss 19.0
DDCD without hidden state loss, $\tau=0.1$ 16.2
DDCD without hidden state loss, $\tau=0.001$ 16.0
DDCD without hidden state loss, $\tau=0$ 15.9
LM baseline 15.2
Table 6: Rouge-L scores for different DDCD model variants.
From these results, we can see that the performance correspondingly converges to that of LM. Experiments are run on the NarrativeQA dataset with phi-3 as the base model. Thus, we conclude that the overall distillation objective is central to the observed gains w.r.t LM.
B.3 Forgetting of General Knowledge after KM training
We believe our approach to be more robust to downsides associated with knowledge injection, given the plug-and-play nature of Knowledge Modules. Indeed, if other information is altered in unintended ways, we always have the opportunity to downweight (or outright remove) the Knowledge Module for a given document. Because we never modify the base model, we can always revert to a previous capability present in the base model. That said, we test how much general knowledge is forgotten after Knowledge Module training, per your suggestion. We evaluate the KMs on the MMLU (Hendrycks et al., 2021) dataset, where for each datapoint, we randomly sample a KM with which to answer the question. We use a subset of 400 examples, evenly split across all MMLU subjects. We compare to the base model, llama 3 8B-Instruct below
Model Accuracy
base model 63.9
base model + KM 62.9
Table 7: Accuracy comparison between the base model and its variant with KM.
We see that the base model equipped with a Knowledge Module is able to retain its overall general knowledge, dropping only 1 point on MMLU.